Grade 11Algebra

1, z1, z2.....z n-1 are the nth roots of unity, then (1-z1)(1-z2)....(1-z n-1) is equal to?

Anikash Chakraborty
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago

ou can see that from

(z-1)(1+z+z2+...+zn-1) = zn - 1 = (z-1)(z-z1)(z-z2)...(z-zk)

Multiplying out shows the left and middle are equal, then the middle and the right have the same (distinct) roots and both have leading coefficient 1, so they're equal too.